A cashier has a total of 33 bills consisting of ones, fives and twenties. The number of twenties is 12 less than the number of ones. The total value of money is $183. How many of each denomination of bill are there?
Thank you.

Answer by Boreal(15235)

x=ones
x-12=twenties.
Their sum is 2x-12, so
33-(2x-12)=fives
x+20(x-12)+5(45-2x)=183
collect terms after distributing
x+20x-240+225-10x=183
11x-15=183
11x=198
x=18
x-12=6
33-(2x-12)=9
There are 33 bills
18 are ones for $18
9 are fives for $45
6 are twenties for $120
Total is 33 for $183.